Question 1 (1 point) TTC VC FC AC AVC AFC MC 0 N 1 2 H 3 2 M 3 G 6 5 What is the correct value for H? 5 2 4 3
Question 2 (1 point) Q TC VC FC AC AVC AFC MC 0 N 1 2 Н 3 2 M 3 G 6 5 What is the correct value for M? 3 5
Question 3 (1 point) Q TC VC FC AC AVC AFC MC 0 N 1 2 Н 3 2 M 3 G 6 5 What is the correct value for G? 3 15 0 2
Question 4 (1 point) Q TC VC FC AC AVC AFC MC O NZ I ܢܙ 3 2 M 3 G 6 5 What is the correct value for N? Оо 03 02 0-3
Answer #1

1. AVC = VC/Q = 2/1 = 2

The correct option is B. H=2.

2. TC for 3 units = AC*Q = 6*3 = 18

VC for 3 units = AVC*Q = 5*3 = 15

FC for all units (since it is same) = 18-15 = 3

TC for 1 unit = 2+3 = 5

TC for 2 units = TC for 1 unit + MC = 5+3 = 8

The correct option is B. M = 8

3. The correct option is A. G = 3.

We have calculated this in previous section.

4. The correct option is A. N = 0.

This is because at 0 unit of output the variable cost is always zero.

NOTE: I HOPE YOU WILL BE SATISFIED WITH MY ANSWER PLEASE DO PROVIDE RATING. THANK YOU AND HAVE A NICE DAY. :))